OpenStudy (anonymous):

Please help!!! I give medalssss c; How do you use trigonometric ratios to solve for a missing side or angle of a right triangle?

OpenStudy (johnweldon1993):

 Let me label the sides and pick an angle...

OpenStudy (johnweldon1993):



OpenStudy (johnweldon1993):

Now...what are our trig functions? \(\large sin \space cos \space tan \space cot \space sec\space csc\)

OpenStudy (anonymous):

The first 3....

OpenStudy (johnweldon1993):

We know that \[\large sin = \frac{opposite}{hypotenuse}\] \[\large cos = \frac{adjacent}{hypotenuse}\] \[\large \tan = \frac{opposite}{adjacent}\] do you know those? or seen them before?

OpenStudy (anonymous):

Yes, i have